OpenStudy (jojo12g):

Which expression has a value of –12? A. –4 – (–8) B. 3 – 9 C. –16 – (–4) D. –6 – 18

OpenStudy (jojo12g):

@M_M13

OpenStudy (alexandervonhumboldt2):

look: A -4-(-8)=-4+8=4 B 3-9=-6 C -16-(-4)=-16+4=-12 D -6-18=-24 -(-)=+

OpenStudy (bluebeta):

C